OpenStudy (anonymous):

FAN AND MEDAL A box has 1 red marble, 1 blue marble, and 4 green marbles of the same shape and size. Dan draws a marble randomly from the box, replaces it, and then draws another marble randomly. What is the probability of drawing two green marbles in a row? Explain your answer.

OpenStudy (anonymous):

Uhh Wouldn't It be 4/6 * 4/6=16/36=4/9

OpenStudy (anonymous):

That is what I think

OpenStudy (anonymous):

but its says to explain the answer to

OpenStudy (anonymous):

Ok